A Carnot engine operating between 527C and 27 C produces 600KW of work. Determine,

(a) Qa, KW; (b) Qr, KW; (c) ΔS during heat rejection, KW/K and (d) thermal efficiency of the cycle.


1
Expert's answer
2022-04-04T09:04:54-0400

Answer

a) relation between temperature and hear are given by

"\frac{Q_a}{Q_r}=\frac{T_a}{T_r}"


b) work done is given

"W=Q_a -Q_r"

c) entropy

"S=\frac{dQ}{T}"

d) efficiency

"\eta=1-\frac{Q_r}{Q_a}"